This is a Cluster 2.
I do homeschooling on Ecot but i need help with the problem
Is (-1,-2) a solution to the system of equations below?
4x + y= -6
2x + 4y= -10

Answer by Alan3354(69443)

Is (-1,-2) a solution to the system of equations below?
4x + y= -6
2x + 4y= -10
-----------
Sub -1 for x and -2 for y and see if the 2 sides are equal.
----
4x + y= -6
4*(-1) + (-2) =? -6
-6 = -6 --> it's a solution of that equation.
--------------
Do the other one the same.